OpenStudy (anonymous):

Show that the integral of (1+x)/(1+x^2)dx = Pi

OpenStudy (anonymous):

\[\int\limits_{-\infty}^{\infty}(1+x)/(1+x^2)dx=\Pi\]

OpenStudy (turingtest):

split up the integral and it should be a snap...

OpenStudy (anonymous):

making it \[\int\limits_{-00}^{oo}1/(1+x^2) + \int\limits_{-00}^{00}x/(1+x^2)\]

OpenStudy (turingtest):

\[\int_{-\infty}^{\infty}\frac{1+x}{1+x^2}dx=\int_{-\infty}^{\infty}\frac1{1+x^2}dx+\int_{-\infty}^{\infty}\frac x{1+x^2}dx\]the first integrand is even and the second is odd, so this becomes\[=2\int_{0}^{\infty}\frac1{1+x^2}+0=2\tan^{-1}x|_{0}^{\infty}=2\lim_{n \rightarrow \infty}\tan^{-1}n=\pi\]badda-bing :D any questions?

OpenStudy (turingtest):

make sense? yes? no? maybe?

OpenStudy (anonymous):

heheh not yet

OpenStudy (anonymous):

how do you identify even and odd is it the power?

OpenStudy (anonymous):

and wheres the 2 comming from?

OpenStudy (turingtest):

an odd function is defined as a function such that\[f(-x)=-f(x)\]and an even function is defined as a function such that\[f(-x)=f(x)\]check to see that the first integrand is even and the second is odd while I type the rest of the explanation

OpenStudy (turingtest):

for all even functions we have the following very useful fact\[\int_{-a}^{a}f(x)=2\int_{0}^{a}f(x)\]think about the area under the graph of x^2 (which is an even function) from x=-a to x=a and this should make sense the area is symmetric !

OpenStudy (turingtest):

for all odd functions we have that\[\int_{-a}^{a}f(x)=0\]this should make sense if you consider the area under x^3 (an odd function) from x=-a to x=a again the areas on each side of x=0 are the same, but they are negative on one side and positive on the other hence the integral is zero

OpenStudy (anonymous):

oh im seeing it

OpenStudy (turingtest):

here are pics of what I'm saying about the areas even: http://www.wolframalpha.com/input/?i=integral+from+-1+to+1+of+x%5E2dx odd: http://www.wolframalpha.com/input/?i=integral+from+-1+to+1+of+x%5E3dx

OpenStudy (zarkon):

the integral does not converge unless you use the Cauchy principal value

OpenStudy (turingtest):

The who-what? what is wrong with my analysis?

OpenStudy (turingtest):

I did see that the integral would not yield pi unless I used the even-odd trick, but I thought I was just messing up...

OpenStudy (anonymous):

cool yea i got it! thanks you!

OpenStudy (turingtest):

but zarkon said something and I want to understand it :/

OpenStudy (zarkon):

when taking the integral \[\int\limits_{-\infty}^{\infty}f(x)dx\] you cant do this \[\lim_{a\to\infty}\int\limits_{-a}^{a}f(x)dx\] ( which is what your odd function trick relies upon)

OpenStudy (turingtest):

so you can't use the odd function thing at all in improper integrals? then how do I solve this?

OpenStudy (zarkon):

you shouldn't

OpenStudy (zarkon):

unless you are finding the Cauchy principal value :)

OpenStudy (turingtest):

I am looking at wikipedia for the Cauchy principle value thing... I don't suppose you'd care to demonstrate it here, eh? I've never heard of it.

OpenStudy (zarkon):

ex \[\lim_{a\to\infty}\int\limits_{-a}^{a}\frac{x}{1+x^2}dx=0\]

OpenStudy (zarkon):

the integral is actually \[\int\limits_{-\infty}^{\infty}\frac{x}{1+x^2}dx=\lim_{a\to-\infty}\lim_{b\to\infty}\int\limits_{a}^{b}\frac{x}{1+x^2}dx\] which does not exist

OpenStudy (turingtest):

the Cauchy principle value is from complex analysis, or what?

OpenStudy (zarkon):

I don't know the original context of how it was first used

OpenStudy (turingtest):

@Lammy Zarkon is saying my proof is flawed, so don't rely on it. Though maybe your teacher will let it slide if they are not as smart as him :P so Zarkon, what is your advice for Lammy and I who obviously don't know this technique yet are we without hope?

OpenStudy (anonymous):

haaha but i liked that way

OpenStudy (zarkon):

if he is supposed to show that the integral is pi then maybe his prof wants him to find the CPV...then your way of doing the integral is fine

OpenStudy (anonymous):

never learn the CPV yet

OpenStudy (zarkon):

though in reality...the integral does not converge

OpenStudy (turingtest):

I can see I need to read about this to know what you are really saying... the integral does not converge, but if I use this CPV then it does...and my trick becomes valid... hm... Until then, I suppose you will have to do with my flawed proof Lammy

OpenStudy (zarkon):

is this for a physics class?

OpenStudy (zarkon):

they tend to like to use the CPV without stating so.

OpenStudy (anonymous):

no its for calulus class

OpenStudy (turingtest):

I'm betting the professor overlooked it because he's not as good as Zarkon ;)

OpenStudy (zarkon):

I would ask your professor about it.

OpenStudy (zarkon):

was the problem from a book or from the prof?

OpenStudy (anonymous):

its from the prof!

OpenStudy (anonymous):

haha yea this prof. makes lots of mistake while lecturing

OpenStudy (anonymous):

i will ask her on monday

OpenStudy (zarkon):

from mathematica v8 Integrate::idiv: Integral of 1/(1+x^2)+x/(1+x^2) does not converge on {-\[Infinity],\[Infinity]}. >> then using the CPV In[6]:= NIntegrate[x/(1 + x^2), {x, -\[Infinity], \[Infinity]}, Method -> {"PrincipalValue"}] Out[6]= 0.

OpenStudy (turingtest):

so it doesn't converge to pi even with the CPV ?

OpenStudy (zarkon):

it will if I did the entire problem

OpenStudy (turingtest):

oh ic, you just did the second integral

OpenStudy (zarkon):

yes...the first integral is ok...the second one is the bad one

OpenStudy (turingtest):

yeah,because at first I didn't see that it was odd, and I got divergence when I evaluated it then I saw the oddness and was like 'aha!', and forgot about trying to explain my earlier issue clearly it is not to be ignored

OpenStudy (zarkon):

just out of curiosity...is your prof a full time prof/adjunct/grad student?

OpenStudy (zarkon):

it should not be ignored :)

OpenStudy (anonymous):

full time prof

OpenStudy (zarkon):

time for a shower....fun talking...later

OpenStudy (anonymous):

thanks a lot

OpenStudy (zarkon):

np

OpenStudy (turingtest):

dinner for me later y'all !